# An interpretation of the Brownian bridge as a physics-informed prior for the Poisson equation

### v1

@@ 4.1 Regularity
The RKHS of eq. (3.4) is the space $H_{k}\coloneqq\{u\in H^{1}(\Omega):u=0,\>\mathrm{on}\>\partial\Omega\}$.
@@
Fix $x_{1}^{\prime},\dots,x_{d}^{\prime}\in\Omega$, with $\Omega=[0,1]^{d}$. If any $x_{1},\dots,x_{d}$ is 00 or $1$, then $k(x_{1},\dots,x_{d},x^{\prime}_{1},\dots,x^{\prime}_{d})=0$. To show $k\in H^{1}(\Omega)$, define the partial sum
$k^{S}(x_{1},\dots,x_{d},x_{1}^{\prime},\dots,x_{d}^{\prime})=2^{d}\sum_{n_{1},\dots n_{d}=1}^{S}\frac{\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})},$
Since the convergence is absolute and uniform, $\exists K>0$ such that for any $(x_{1},\dots,x_{d})\in\Omega$,
$2^{d}\sum_{n_{1},\dots,n_{d}\in\mathbb{N}}\frac{|\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})|}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}\leq K.$
Now for $x=(x_{1},\dots,x_{d})$ and $y=(y_{1},\dots,y_{d})$ in $\Omega$,
$|k^{S}(x_{1},\dots,x_{d},x_{1}^{\prime},\dots,x_{d}^{\prime})-k^{S}(y_{1},\dots,y_{d},x_{1}^{\prime},\dots,x_{d}^{\prime})|=$
$\left|2^{d}\sum_{n_{1},\dots n_{d}=1}^{S}\frac{\left[\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})-\sin(n_{1}\pi y_{1})\cdots\sin(n_{d}\pi y_{d})\right]\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}\right|$
$\leq 2^{d}\sum_{n_{1},\dots n_{d}=1}^{S}\left|\frac{\left[\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})-\sin(n_{1}\pi y_{1})\cdots\sin(n_{d}\pi y_{d})\right]\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}\right|$
$=2^{d}\sum_{n_{1},\dots n_{d}=1}^{S}\frac{\left|\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})-\sin(n_{1}\pi y_{1})\cdots\sin(n_{d}\pi y_{d})\right||\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})|}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}$
$\leq 2^{d}\sum_{n_{1},\dots n_{d}=1}^{S}\frac{|\sin(n_{1}\pi x^{\prime}_{1})\cdots\sin(n_{1}\pi x^{\prime}_{d})|}{\pi^{2}(n_{1}^{2}+\cdots+n_{d}^{2})}\|x-y\|\leq K\|x-y\|,$
where we have used the fact that $\sin(n_{1}\pi x_{1})\cdots\sin(n_{d}\pi x_{d})$ is Lipschitz with constant $1$. As the Lipschitz constant of $k^{S}$ is bounded for any $S$, and $\lim_{S\to\infty}k^{S}=k$ uniformly, $k$ is also Lipschitz continuous. Hence, $k$ is weakly differentiable. We have shown that $k\in H_{k}$, satisfying property (i) of Definition 2.1.
$\langle u,k(\cdot,x^{\prime})\rangle_{H_{k}}=\sum_{\alpha\in\mathbb{N}^{d}}\lambda_{\alpha}^{-1}\langle u,\psi_{\alpha}\rangle\langle k(\cdot,x^{\prime}),\psi_{\alpha}\rangle,$
where $\psi_{\alpha}$ is any orthonormal basis. Here, we have used the multi-index notation $\alpha=(n_{1},\dots,n_{d})$. Pick the basis to be the d-dimensional Fourier sine series $\psi_{\alpha}=2^{d/2}\sin(\alpha\pi x)$. We can expand $u$ by $u=\sum_{\alpha\in\mathbb{N}^{d}}\langle u,\psi_{\alpha}\rangle\psi_{\alpha}$. Now, we have (by Lebesgue’s dominated convergence theorem) for any fixed $\alpha$
[…]
$=\int\sum_{\gamma\in\mathbb{N}^{d}}\left\{2^{3d/2}\lambda_{\gamma}\sin(\gamma\pi x)\sin(\alpha\pi x)\sin(\gamma\pi x^{\prime})\right\}dx$
[…]
$=2^{d/2}\lambda_{\alpha}\sin(\alpha\pi x^{\prime}),$
where the last line holds as $\langle 2^{d}\sin(\gamma\pi x),\sin(\alpha\pi x)\rangle=1$ for $\alpha=\gamma$ and 00 otherwise (it is the orthonormal basis we picked).
$\langle u,k(\cdot,x^{\prime})\rangle_{H_{k}}$ $=\sum_{\alpha\in\mathbb{N}^{d}}\lambda_{\alpha}^{-1}\langle u,\psi_{\alpha}\rangle\langle k(\cdot,x^{\prime}),\psi_{\alpha}\rangle$
$=\sum_{\alpha\in\mathbb{N}^{d}}\lambda_{\alpha}^{-1}\langle u,\psi_{\alpha}\rangle 2^{d/2}\lambda_{\alpha}\sin(\alpha\pi x^{\prime})$
$=\sum_{\alpha\in\mathbb{N}^{d}}\langle u,\psi_{\alpha}\rangle 2^{d/2}\sin(\alpha\pi x^{\prime})$
$=\sum_{\alpha\in\mathbb{N}^{d}}\langle u,\psi_{\alpha}\rangle\psi_{\alpha}(x^{\prime})=u(x^{\prime}).$
This shows requirement (ii) of Definition 2.1 also holds, and $k$ is the unique reproducing kernel for $H_{k}$. ∎
The above result provides us with an interesting method to prove the well-known result that Brownian bridge sample paths are nowhere differentiable. This follows immediately by combining Lemma 4.1 and Theorem 2.2.
@@
Definition 4.1 (Version of a stochastic process [11]).
@@
We will look to find versions of the Brownian bridge on powers of its RKHS:
Definition 4.2 (Powers of RKHS [43]).
@@
Note we have the property $H_{k}=H_{k}^{1}\subset H_{k}^{p_{1}}\subset H_{k}^{p_{2}}\subset L^{2}(\Omega)$, for all $10<p_{2}<p_{1}<$ 1. Evidently, as $p$ decreases, the power RKHS loses some regularity. Note that $H_{k}^{p}$ is itself a RKHS with kernel $k^{p}$. Finally, we need the following theorem, which follows from Driscol’s theorem [18, Theorem 3].
Theorem 4.1 (Theorem 4.12 [29]).
@@
Let $u$ be the unit Brownian bridge with $d=1$. Then, for all $11/2<p<$ 1, there exists a version of $u$, $\tilde{u}$, such that $\tilde{u}\in H_{k}^{p}$ with probability one.
@@
First, we need to check when the condition $\sum_{n\in\mathbb{N}}\lambda_{n}^{p}\psi_{n}^{2}(x)<\infty$ for all $x\in J$ holds. The eigenvalues and eigenfunctions are $(n^{2}\pi^{2})^{-1}$ and $\phi_{n}(x)=\sqrt{2}\sin(n\pi x)$, $n\in\mathbb{N}$, respectively. Then for any $x\in J$,
@@
when $1/2<p\leq 1$, which can be verified by the $p$-series test. We now will show (i) holds. We have $\sum_{n\in\mathbb{N}}\frac{1}{(n^{2}\pi^{2})^{1-p}}<\infty$ for any $11/2<p<$ 1, which proves the result. ∎
The proposition shows that we can find a version of the Brownian bridge which is, in a sense, as close as possible to being an $H^{1}(\Omega)$ function without being weakly differentiable. If desired, one can construct these versions using the Karhunen-Loeve expansion (KLE) of the $p$-power kernel.
While at first the poor regularity of the prior may feel discouraging, the fact that the sample paths a.s. do not belong to the solution space of the Poisson equation is less of an issue than seems. Recall by Proposition 2.1 that the posterior mean function $\tilde{m}$ will live in the RKHS of the prior covariance. As a result, the regularity of $\tilde{m}$ will match the desired behavior required of the energy functional, i.e., an $H^{1}(\Omega)$ function which satisfies the boundary conditions. In regression tasks we often interpret $\tilde{m}$ to be the predictor, with the variance representing a worst-case error, so in this sense, it is ideal that the RKHS is a first-order Sobolev space. In fact, the RKHS being norm-equivalent to a Sobolev space is a crucial hypothesis needed to establish convergence conditions, explored later.
@@ 4.2 Finite-dimensional representations
We must work with a finite-dimensional representation of the prior in practical applications. In most uses of GP regression a mesh of test points is placed on $\Omega$ where the posterior predictions are queried. Instead, we derive a finite-dimensional basis approximation to the prior which places the mesh on $L^{2}(\Omega)$. We begin with the 1D case for demonstration. Without loss of generality, we derive the results with $\beta=1$. Since we do not permit $\beta$ to be zero or infinite, the results do not change for different values of $\beta$.
Formally, by assuming the existence of the Lebesgue measure on $L^{2}(\Omega)$, we write the physics-informed prior $\mu_{C}\sim\mathcal{N}(0,C)$ as
$\mu_{C}(du)=\frac{1}{Z}\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}\mathcal{D}u,$ (4.1)
where we have centered the measure, and $Z$ is the normalization constant. The shift is justified since $u^{0}\in H_{p}\subset H_{k}$, which results in an equivalent measure. Here, $\mathcal{D}u$ serves as a replacement for the non-existent Lebesgue measure in infinite-dimensions. This idea appears in different path integral approaches for Bayesian inverse problems, including Bayesian field theory [33], information field theory [21], physics-informed information field theory [3, 25], and others [14]. Of course, eq. (4.1) is not well-defined in the continuum limit, but, as the physicists do, we will look to extract meaning from this expression. The reference [24] provides a mathematical background to the nuances of using such definitions.
@@
Pick any orthonormal basis [^1] One could also choose a grid of piecewise constant functions on $\Omega$, which corresponds to picking test points. in $L^{2}(\Omega)$, $(\psi_{i})_{\in\mathbb{N}}$, and let $\mathcal{F}\subset L^{2}(\Omega)$ be the set =$\mathcal{F}=\{u\in L^{2}(\Omega):u=\sum_{i=1}^{n}\alpha_{i}\psi_{i}\}$ for fixed $n\in\mathbb{N}$. Then, $\dim(\mathcal{F})=n<\infty$. Let the restriction of $C$ to $\mathcal{F}$ be given by $\Sigma_{\mathcal{F}}$. In this case, $\Sigma_{\mathcal{F}}$ is the covariance matrix of a unique finite-dimensional Gaussian measure appearing as
$\mu^{(n)}(d\hat{u})=\frac{1}{\sqrt{(2\pi)^{n}|\Sigma_{\mathcal{F}}|}}\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}}^{-1}\hat{u}\rangle\right\}d\hat{u},$ (4.2)
where the Lebesgue measure induced by the $L^{2}(\Omega)$-inner product, $d\hat{u}$, is well-defined. So, eq. (4.2) can be regarded as a measure over a finite-dimensional space on the cylinder sets of $\mathcal{F}$. The next series of results show that this measure has the correct limiting behavior.
@@
Let $\mathcal{F}_{1}\subset\mathcal{F}_{2}\subset L^{2}(\Omega)$ with $\dim(\mathcal{F}_{1})=n_{1}\leq\dim(\mathcal{F}_{2})=n_{2}<\infty$ and $C$ be a covariance operator on $L^{2}(\Omega)$. Now, let the restriction of $C$ to $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ be given by $\Sigma_{\mathcal{F}_{1}}$ and $\Sigma_{\mathcal{F}_{2}}$, respectively. Then, both $\Sigma_{\mathcal{F}_{1}}$ and $\Sigma_{\mathcal{F}_{2}}$ uniquely define the finite-dimensional Gaussian measures
$\mu^{(n_{1})}(d\hat{u})$ $=\frac{1}{\sqrt{(2\pi)^{n_{1}}|\Sigma_{\mathcal{F}_{1}}|}}\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}_{1}}^{-1}\hat{u}\rangle\right\}d\hat{u}$
$\mu^{(n_{2})}(d\hat{u})$ $=\frac{1}{\sqrt{(2\pi)^{n_{2}}|\Sigma_{\mathcal{F}_{2}}|}}\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}_{2}}^{-1}\hat{u}\rangle\right\}d\hat{u},$
respectively on $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ cylinder sets, where $d\hat{u}$ is the Lebesgue measure. In addition to this, the restriction of $\mu^{(n_{2})}$ to $\mathcal{F}_{1}$ cylinder sets is exactly $\mu^{(n_{1})}$.
@@
Note the $\mathcal{F}_{1}$ cylinder sets are also cylinder sets in $\mathcal{F}_{2}$. Both $\mu_{C}$ and $\mu^{(n_{2})}$, when restricted to $\mathcal{F}_{1}$ cylinder sets, define Gaussian measures on $\mathcal{F}_{1}$, uniquely determined by their covariances. The measure $\mu^{(n_{1})}$ has covariance $\Sigma_{\mathcal{F}_{1}}$ by definition, while restriction of $\mu^{(n_{2})}$ to $\mathcal{F}_{1}$ cylinder sets has covariance $\Sigma_{\mathcal{F}_{2}}\restriction_{\mathcal{F}_{1}}=\Sigma_{\mathcal{F}_{1}}$. ∎
The intuition behind the result is as follows. In applications, we place a mesh, described by $\mathcal{F}_{1}$, on $L^{2}(\Omega)$ so that we are in the finite-dimensional setting for sampling. We pick this mesh by truncating the chosen orthonormal basis at some point. Refining the mesh further to $\mathcal{F}_{2}$ does not change how the prior behaves on the original mesh because $\mu^{(n_{1})}$ and $\mu^{(n_{2})}$ agree on the $\mathcal{F}_{1}$ cylinder sets. Practically this means that in applications there is some cutoff point where refining the mesh any further does not reasonably change the results.
Lemma 4.2 shows that $\mu^{(n)}$ is finitely additive, so that it is indeed a well-defined measure. We also have that $\mu^{(n)}$ is equivalent to the Lebesgue measure when restricted to a finite-dimensional space $\mathcal{F}$, hence it is regular. Putting these facts together, we can prove countable additivity on $L^{2}(\Omega)$:
Proposition 4.3.
Let $\mu$ be a finitely additive regular measure defined on $\sigma(\mathcal{R})$. Then $\mu$ is also countably additive on $L^{2}(\Omega)$.
Proof.
Recall that $\sigma(\mathcal{R})=\mathcal{B}(L^{2}(\Omega))$, so it is sufficient to prove the result for $\sigma(\mathcal{R})$. Take $I=\cup_{n=1}^{\infty}I_{n}$ to be a disjoint union of cylinder sets. Then, $I$ is a cylinder set. Let $I_{0}=\mathcal{R}-I$. For countable additivity, we must show that $\sum_{n=0}^{\infty}\mu(I_{n})=1$. By finite additivity of $\mu$, we have
$\sum_{n=0}^{\infty}\mu(I_{n})=\lim_{R\to\infty}\sum_{n=0}^{R}\mu(I_{n})=\lim_{R\to\infty}\mu\left(\cup_{n=0}^{R}I_{n}\right)\leq 1.$
To show the reverse, we will use the fact that $\mu$ is regular and it suffices to show that $\sum_{n=0}^{\infty}\mu(G_{n})\geq 1$ with each $G_{n}$ an open cylinder set with $I_{n}\subset G_{n}$. Fix $\varepsilon,r>0$, and let $B(r)$ be a ball in $L^{2}(\Omega)$. Since $B(r)$ is weakly compact, there exists a finite number of sets $G_{0},\dots,G_{R}$ which form an open cover of $B(r)$, and without loss of generality, we can take each $G_{n}$, $n=0,\dots,R$ to be an open cylinder set.
Now, define $G=\mathcal{R}-\cup_{n=0}^{R}G_{n}$. Then $G$ is also an open cylinder set which is disjoint from $B(r)$. Hence,
$\varepsilon\geq\mu(G)\geq 1-\sum_{n=0}^{R}\mu(G_{n})\geq 1-\sum_{n=0}^{\infty}\mu(G_{n}).$
Therefore, $\sum_{n=0}^{\infty}\mu(G_{)}\geq 1-\varepsilon$, and by regularity, $\sum_{n=0}^{\infty}\mu(I_{n})\geq 1-\varepsilon$. Taking the infimum as $\varepsilon\to 0$, we have $\sum_{n=0}^{\infty}\mu(I_{n})\geq 1$, which proves the result. ∎
We now show eq. (4.2) converges to the correct measure in the limit.
@@
Let $d=1$, $(\psi_{i})_{i\in\mathbb{N}}$ be an orthonormal basis for $L^{2}(\Omega)$, and for each $n\in\mathbb{N}$, let $\mu^{(n)}$ be given by eq. (4.2). Then $\mu^{(n)}\implies\mu_{C}$. That is, the sequence $(\mu^{(n)})_{n\in\mathbb{N}}$ converges weakly to the Gaussian measure $\mu_{C}=\mathcal{N}(0,C)$ on $L^{2}(\Omega)$.
@@
$\phi_{\mu^{(n)}}(u)=\exp\left\{-\frac{1}{2}\left\langle\sum_{i=1}^{n}\alpha_{i}\psi_{i},\Sigma_{\mathcal{F}}\left(\sum_{i=1}^{n}\alpha_{i}\psi_{i}\right)\right\rangle\right\},$
where $\alpha_{i}=\langle u,\psi_{i}\rangle$. We have $\lim_{n\to\infty}\phi_{\mu^{(n)}}(u)=\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}$, which is the characteristic function of $\mathcal{N}(0,C)$ [32, Lemma 2.1]. Convergence in characteristic functions implies weak convergence in measure [9]. ∎
While Theorem 4.2 is valid only in 1D (so that the Gaussian measure is well-defined), we can provide a similar convergence condition in the setting of Lévy white noises (generalized random fields) for $d>1$, provided that we let $\Omega=\mathbb{R}^{d}$. Recall the Schwartz space of smooth, rapidly decaying functions
Let $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ denote the dual of $\mathcal{S}(\mathbb{R}^{d})$. Note that $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is known as the space of tempered distributions, due to the test function topology of $\mathcal{S}\left(\mathbb{R}^{d}\right)$. We can rely on Lévy’s continuity theorem [8] in the setting of tempered distributions:
Theorem 4.3 (Lévy’s continuity theorem [8]).
Let $(\mu^{(n)})_{n\in\mathbb{N}}$ be a sequence of Lévy white noises, each with characteristic function $\phi_{\mu^{(n)}}$. Suppose $\phi_{\mu^{(n)}}$ converges pointwise to some function $\phi:\mathcal{S}\to\mathbb{C}$, which is continuous at 00. Then $\phi$ is the characteristic function of some Lévy white noise $\mu$ on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Further, $\mu^{(n)}$ converges in distribution to $\mu$.
This yields a parallel convergence theorem to Theorem 4.2 for the general case.

### v2

@@ 4.1 Regularity
The RKHS of eq. (3.4) is the space $H_{k}\coloneqq H^{1}_{0}(\Omega)=\{u\in H^{1}(\Omega):u=0,\>\mathrm{on}\>\partial\Omega\}$.
@@
The proof is in Appendix B. ∎
The above result provides us with an interesting method to prove the well-known result that Brownian bridge sample paths are nowhere differentiable. This follows immediately by combining Lemma 4.1 and Proposition 2.2.
@@
Definition 4.1 (Version of a stochastic process [bremaud2014fourier]).
@@
We will look to find versions of the Brownian bridge on powers of its RKHS, also known as a Hilbert scale [mathe2008direct]:
Definition 4.2 (Powers of RKHS [steinwart2008support]).
@@
Note we have the property $H_{k}=H_{k}^{1}\subset H_{k}^{p_{1}}\subset H_{k}^{p_{2}}\subset L^{2}(\Omega)$, for all $0<p_{2}<p_{1}<1$. Evidently, as $p$ decreases, the power RKHS loses some regularity. Note that $H_{k}^{p}$ is itself a RKHS with kernel $k^{p}$. Finally, we need the following theorem, which follows from Driscol’s theorem [driscoll1973reproducing, Theorem 3].
Theorem 4.1 (Theorem 4.12 [kanagawa2018gaussian]).
@@
Let $u$ be the unit Brownian bridge with $d=1$. Then, for all $1/2<p<1$, there exists a version of $u$, $\tilde{u}$, such that $\tilde{u}\in H_{k}^{p}$ with probability one.
@@
First, we need to check when the condition $\sum_{n\in\mathbb{N}}\lambda_{n}^{p}\psi_{n}^{2}(x)<\infty$ for all $x\in\Omega$ holds. The eigenvalues and eigenfunctions are $(n^{2}\pi^{2})^{-1}$ and $\phi_{n}(x)=\sqrt{2}\sin(n\pi x)$, $n\in\mathbb{N}$, respectively. Then for any $x\in\Omega$,
@@
when $1/2<p\leq 1$, which can be verified by the $p$-series test. We now will show (i) holds. We have $\sum_{n\in\mathbb{N}}\frac{1}{(n^{2}\pi^{2})^{1-p}}<\infty$ for any $1/2<p<1$, which proves the result. ∎
The proposition shows that we can find a version of the Brownian bridge which is, in a sense, as close as possible to being an $H_{0}^{1}(\Omega)$ function without being weakly differentiable. If desired, one can construct these versions using the Karhunen-Loève expansion (KLE) of the $p$-power kernel.
While at first the poor regularity of the prior may feel discouraging, the fact that the sample paths a.s. do not belong to the solution space of the Poisson equation is less of an issue than seems. In regressions tasks we often interpret $\tilde{m}$ to be the predictor, with the variance representing a worst-case error, and the individual sample paths are inconsequential. Recall by Proposition 2.1 that the posterior mean function $\tilde{m}$ will live in the RKHS of the prior covariance. As a result, the regularity of $\tilde{m}$ will match the desired behavior required of the energy functional, i.e., an $H_{0}^{1}(\Omega)$ function, so in this sense, it is ideal that the RKHS is a first-order Sobolev space. We do not add any additional smoothness assumptions beyond what is needed for the PDE solution to exist. In fact, the RKHS being norm-equivalent to a Sobolev space is a crucial hypothesis needed to establish convergence conditions, explored later.
@@ 4.2 Finite-dimensional representations
We must work with a finite-dimensional representation of the prior in practical applications. In most uses of GP regression a pointwise mesh of test points is placed on $\Omega$ where the posterior predictions are queried. Instead, we derive a finite-dimensional basis approximation to the prior in $L^{2}(\Omega)$. The motivation behind this is to enable scalable prior/posterior approximation algorithms by placing the computational burden on sampling basis function coefficients, rather than on the test points. We prove convergence in measure and in Wasserstein distance under our approximation. Without loss of generality, we derive the results with $\beta=1$. Since we do not permit $\beta$ to be zero or infinite, the results do not change for different values of $\beta$.
Our finite-dimensional representation is inspired by the mathematics behind quantum field theory, where Gaussian measures are oftentimes expressed under the path integral formalism. Formally, by assuming the existence of the Lebesgue measure on $L^{2}(\Omega)$, we write the physics-informed prior $\mu_{C}\sim\mathcal{N}(0,C)$ as
$\mu_{C}(du)``="\frac{1}{Z}\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}\mathcal{D}u,$ (4.1)
where we have centered the measure, and $Z=\int_{L^{2}(\Omega)}\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}\mathcal{D}u$ is the normalization constant. The shift is justified since $u^{0}\in H_{k}$, which results in an equivalent measure. Here, $\mathcal{D}u$ serves as a replacement for the non-existent Lebesgue measure in infinite-dimensions. This idea appears in different path integral approaches for Bayesian inverse problems, including Bayesian field theory [lemm2003bayesian], information field theory [ensslin2009information], physics-informed information field theory [alberts2023physics, hao2024information], and others [chang2014path]. Of course, eq. (4.1) is not well-defined in the continuum limit, but, as the physicists do, we will look to extract meaning from this expression. The reference [glimm2012quantum] provides a mathematical background to the nuances of using such definitions.
@@
Pick any orthonormal basis [^1] One could choose a grid of piecewise constant functions on $\Omega$, which corresponds to picking test points. in $L^{2}(\Omega)$, $(\psi_{i})_{i\in\mathbb{N}}$, and let $\mathcal{F}_{n}\subset L^{2}(\Omega)$ be the set $\mathcal{F}_{n}=\{u\in L^{2}(\Omega):u=\sum_{i=1}^{n}a_{i}\psi_{i},a_{i}\in\mathbb{R}\}$ for fixed $n\in\mathbb{N}$. Then, $\dim(\mathcal{F}_{n})=n<\infty$. Let $\Sigma_{\mathcal{F}_{n}}$ be the restriction-corestriction of $C$ by both domain and codomain to $\mathcal{F}_{n}$. In this case, $\Sigma_{\mathcal{F}_{n}}$ is the covariance matrix of a unique finite-dimensional Gaussian measure appearing as
$\mu_{n}(d\hat{u})=\frac{1}{\sqrt{(2\pi)^{n}|\Sigma_{\mathcal{F}_{n}}|}}\exp\left\{-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}_{n}}^{-1}\hat{u}\rangle\right\}d\hat{u}$, (4.2)
where the Lebesgue measure induced by the $L^{2}(\Omega)$-inner product, $d\hat{u}$, is well-defined. So, eq. (4.2) can be regarded as a measure over a finite-dimensional space on the cylinder sets of $\mathcal{F}_{n}$. The next series of results show that this measure has the correct limiting behavior.
@@
Let $\mathcal{F}_{n}\subset\mathcal{F}_{m}\subset L^{2}(\Omega)$ with $\dim(\mathcal{F}_{n})=n\leq\dim(\mathcal{F}_{m})=m<\infty$ and $C$ be a covariance operator on $L^{2}(\Omega)$. Let the restriction-corestriction of $C$ to $\mathcal{F}_{n}$ and $\mathcal{F}_{m}$ be given by $\Sigma_{\mathcal{F}_{n}}$ and $\Sigma_{\mathcal{F}_{m}}$, respectively, and let $\mu_{n}$ and $\mu_{m}$ be as given in eq. (4.2) for each. Then, the restriction-corestriction of $\mu_{m}$ to $\mathcal{F}_{n}$ cylinder sets is exactly $\mu_{n}$.
@@
Note the $\mathcal{F}_{n}$ cylinder sets are also cylinder sets in $\mathcal{F}_{m}$. Both $\mu_{C}$ and $\mu_{m}$, when restricted to $\mathcal{F}_{n}$ cylinder sets, define Gaussian measures on $\mathcal{F}_{n}$, uniquely determined by their covariances. The measure $\mu_{n}$ has covariance $\Sigma_{\mathcal{F}_{n}}$ by definition, while restriction of $\mu_{m}$ to $\mathcal{F}_{n}$ cylinder sets has covariance given by the restriction-corestriction of $\Sigma_{\mathcal{F}_{m}}$ to $\mathcal{F}_{n}$, which is $\Sigma_{\mathcal{F}_{n}}$. ∎
The intuition behind the result is as follows. In applications, we represent the prior as a truncated expansion corresponding to a set $\mathcal{F}_{n}\subset L^{2}(\Omega)$ so that we are in the finite-dimensional setting for sampling. Adding additional terms, which corresponds to approximation in $\mathcal{F}_{m},$ does not change how the prior behaves on $\mathcal{F}_{n}$, as $\mu_{n}$ and $\mu_{m}$ agree on the $\mathcal{F}_{n}$ cylinder sets. Practically this means that in applications there is some cutoff point where refining the approximation any further does not reasonably change the results.
Once we have selected a finite-dimensional representation, a straightforward application of Bayes’ rule reveals the posterior. To keep the notation consistent with the infinite-dimensional setting, the resulting posterior, denoted $\mu^{y}_{n}$ can be described by the Radon-Nikodym derivative following Theorem 3.3. That is, after describing the measurement process as a potential $\Phi$, we write
$\frac{d\mu^{y}_{n}}{d\mu_{n}}(\hat{u})\propto\exp\{-\Phi(\hat{u})\},\quad\hat{u}\in\mathcal{F}_{n}.$ (4.3)
In what follows, we show that the finite-dimensional representation of the prior and the resulting posterior have the correct convergence behavior. We begin with the 1D case, and later study the multidimensional case, which is much more delicate.
The first result is on the convergence of the prior.
@@
Let $d=1$, $(\psi_{i})_{i\in\mathbb{N}}$ be an orthonormal basis for $L^{2}(\Omega)$, and for each $n\in\mathbb{N}$, let $\mu_{n}$ be given by eq. (4.2). Then $\mu_{n}\implies\mu_{C}$. That is, the sequence $(\mu_{n})_{n\in\mathbb{N}}$ converges weakly to the Gaussian measure $\mu_{C}=\mathcal{N}(0,C)$ on $L^{2}(\Omega)$.
@@
$\phi_{\mu_{n}}(u)=\exp\left\{-\frac{1}{2}\left\langle\sum_{i=1}^{n}a_{i}\psi_{i},\Sigma_{\mathcal{F}}\left(\sum_{i=1}^{n}a_{i}\psi_{i}\right)\right\rangle\right\},$
where $a_{i}=\langle u,\psi_{i}\rangle$. We have $\lim_{n\to\infty}\phi_{\mu_{n}}(u)=\exp\left\{-\frac{1}{2}\langle u,Cu\rangle\right\}$, which is the characteristic function of $\mathcal{N}(0,C)$ [kuo2006gaussian, Lemma 2.1]. Convergence in characteristic functions implies weak convergence in measure [billingsley2017probability]. ∎
It remains to show that posterior convergence also holds. For this, we rely on [sprungk2020local], which provides the result in terms of Wasserstein distance between measures. That is, under some light assumptions about the potential $\Phi$, if $W_{p}(\mu_{C},\mu_{n})\to 0$ it follows that $W_{p}(\mu_{C}^{y},\mu_{n}^{y})\to 0$ for a given $p\geq 1$. What we have shown thus far is that the sequence of priors $(\mu_{n})_{n\in\mathbb{N}}$ converges weakly in measure to $\mu_{C}$. However, it is not immediately obvious that this leads to $W_{p}$ convergence. It turns out that this does hold provided that the $p$-th moments of $\mu_{n}$ also converge to the $p$-th moment of $\mu_{C}$, see [villani2008optimal, Theorem 6.9]. This leads us to the following lemma.
Lemma 4.2.
Let the conditions of Theorem 4.2 hold, $X_{n}\sim\mu_{n}$, and $X\sim\mu_{C}$. Then for every $p\geq 1$, in the limit $n\to\infty,$ we have $\mathbb{E}[\|X_{n}\|^{p}]\to\mathbb{E}[\|X\|^{p}]$.
Proof.
Let $p\geq 1$ be fixed. Our choice of a truncated basis defines a projection mapping from $L^{2}(\Omega)$ onto $\mathcal{F}_{n}$ for any $n$, which we call $P_{n}$. Then, for every $n$, the random variable $X_{n}$ is given by $X_{n}=P_{n}X$. We will show that $\|X_{n}\|^{p}\to\|X\|^{p}$, which will give the desired moment convergence.
To start, note that $P_{n}\to I$ (the identity operator on $L^{2}(\Omega)$), so for all $u\in L^{2}(\Omega)$, we have $P_{n}u\to u$. Application with the random element $X$ yields
$\|X_{n}-X\|=\|P_{n}X-X\|\to 0,\quad\text{a.s.},$
and by continuity of the norm, this gives $\|X_{n}\|^{p}\to\|X\|^{p}$ with probability $1$. Also, since $\|P_{n}u\|\leq\|u\|$ $\forall u\in L^{2}(\Omega)$, it follows that $\|X_{n}\|^{p}\leq\|X\|^{p}$ a.s. for any $n$.
Now, as $X$ is Gaussian, the Fernique theorem applies, which states that for some $\alpha>0$, $\mathbb{E}[\exp(\alpha\|X\|^{2})]<\infty$ [stuart2010inverse, Theorem 6.9]. In particular, this shows that $\mathbb{E}[\|X\|^{p}]<\infty$. Hence, by the dominated convergence theorem, we find
$\mathbb{E}[\|X_{n}\|^{p}]\to\mathbb{E}[\|X\|^{p}],$
which shows the desired convergence in moments. ∎
We can now prove convergence of the posterior approximation.